OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. Type your answer in the box without skipping any spaces. m = 7, point (-2, 0)

OpenStudy (anonymous):

there are 2 formula you can solve 1) y-y1= m(x-x1) or : y=mx+b